OpenStudy (anonymous):

Tiffany, Lori, and Mika are practicing for an egg-tossing contest. The distance from Tiffany to Lori is 17 inches. The distance from Lori to Mika is 32 inches. The distance from Mika to Tiffany is 28 inches. Find the measures of the three angles in the triangle.

OpenStudy (anonymous):

Law of cosines!

OpenStudy (anonymous):

\[c^2 = a^2 + b^2 - 2ab\cos\theta\]